OpenStudy (anonymous):

please help i dont understand! factor completely. If a polynomial is prime, state this. -12x^2-28x+24

OpenStudy (anonymous):

All the coefficients are even numbers; you can start there.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Factoring out a -1 might also make life easier, but isn't essential..

OpenStudy (anonymous):

i know that the -12is taken out like this -12(but cant figure out this part)

OpenStudy (anonymous):

hmmm, if you do that, then you'll get a fraction, which will make it more difficult.

OpenStudy (anonymous):

(28 is not a multiple of 12)

OpenStudy (anonymous):

does this make it a prime polynomial

OpenStudy (anonymous):

Nope, it's factorable.

OpenStudy (anonymous):

ok but now i am completly lost after reaing the books I thought that is how you do it

OpenStudy (anonymous):

Find the GCF of -12, -28, and 24; take that out and you'll be left with a quadratic trinomial that is the product of two linear binomials.

OpenStudy (anonymous):

4 i believe

OpenStudy (anonymous):

Right, and again, this is just a personal preference because I like my leading coefficients to be positive, but I would take out a -4.

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

\[-12x^2-28x+24=-4(3x^2+7x-6)\] So keep the -4 on the outside and try to factor the 3x^2+7x-6. My preferred way is to multiply the leading coefficient (3) by the constant (-6), then find factors of that number that add up to the middle coefficient (7).

OpenStudy (anonymous):

That way, you can split the middle term and do factor-by-grouping.

OpenStudy (anonymous):

sorry writting this down also to understand it

OpenStudy (anonymous):

There are some guess-and-check shortcuts that are sometimes quicker, but this is the general way that works all the time.

OpenStudy (anonymous):

Please do. Take your time.

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

This is a vital skill for algebra. Really understanding this will make a lot of other things a lot easier.

OpenStudy (anonymous):

6 *1 =6 but if you do 6+1 =7 so would it be (3x+6)(x+1)

OpenStudy (anonymous):

Don't forget that 3 in front of the x^2. (3x+6)(x+1)=3x^2+9x+6, which is not the trinomial you started with.

OpenStudy (anonymous):

nope that does not work when you use the foil system

OpenStudy (anonymous):

3 × -6 = -18 Find factors of -18 that add to 7.

OpenStudy (anonymous):

9*-2=-18 and 9-2 =7 if that is what you are asking me.

OpenStudy (anonymous):

it doesnt work when you foil it using these numbers

OpenStudy (anonymous):

do i have the right idea of what I am doing

OpenStudy (anonymous):

There is another factoring step, though. \[3x^2+7x-6=3x^2+9x-2x-6=(3x^2+9x)+(-2x-6)\] You can factor those two expressions in parentheses separately and find your binomial factors.

OpenStudy (anonymous):

You do have the right idea, but there is just a little more breaking apart left to do.

OpenStudy (anonymous):

i am not sure what you mean by factoring these two expressions

OpenStudy (anonymous):

factor 3x^2+9x and then factor -2x-6

OpenStudy (anonymous):

You'll see that both of them have a common binomial factor of (x+3).

OpenStudy (anonymous):

ok working on it will be back in a min doing it on paper

OpenStudy (anonymous):

In the meantime, here's the guess-and-check method: To factor 3x^2+7x-6, you know that to get the 3x^2 you need to multiply (3x)(x), and to make the -6, you'll need (-1)(6), (1)(-6), (2)(-3), or (-2)(3). You can test those possibilities by FOILing (3x + __)(x + __)

OpenStudy (anonymous):

is it (3x-3)(x+3) and (-2+2)(x+3) still not sure that is right

OpenStudy (anonymous):

Almost . . . \[3x^2+9x=3x(x+3), \space -2x-6=-2(x+3)\] Factor out the (x+3) and you'll get (x+3)(3x-2).

OpenStudy (anonymous):

Don't forget to include the -4 that got factored out at the beginning and you're done.

OpenStudy (anonymous):

oh i was doing it wrong ty

OpenStudy (anonymous):

how do you write that

OpenStudy (anonymous):

ty for your help and being patient with me

OpenStudy (anonymous):

And as a taste of things to come, if you had the function y=-12x^2-28x+24, the X-intercepts of the graph will be at -3 and 2/3 I'll let you figure out why.

OpenStudy (anonymous):

lol ok i will try to figure that out